A car does 63062.0 J of work in travelling 3.0 km at a constant speed. What was the average retarding force from all sources acting on the car ?
Answer #1

the statement constant speed (and I assume, in a straight line), tells you that the net force acting on the car is zero (otherwise the velocity would vary if net forces produced accelerations)

this means that the force generated by the car was equalled by the sum of all retarding forces;

now, use W= F d => F=W/d=63452J/2600m = 24.4N

since the forces opposing the car must equal the force generated by the car, we know that the sum of all retarding forces is -24.4N
Answer #2

W = F*S

63062 = F * 3000

retarding force, F = 21.021 N
Answer #3

the statement constant speed (and I assume, in a straight line), tells you that the net force acting on the car is zero (otherwise the velocity would vary if net forces produced accelerations)

this means that the force generated by the car was equalled by the sum of all retarding forces;

now, use W= F d => F=W/d=63062J/3000m = 21.02N

since the forces opposing the car must equal the force generated by the car, we know that the sum of all retarding forces is -21.02N